1. A survey of a lake revealed that there are 28 bass living in the water. If 28 is 16% of the total fish population in the lake, then how many total fish are in the lake?
Answer by Boreal(15235)

28 bass are 16%.
this is about 1/6, so there should be about 170 total, which is just over 6 times 16%
Exact answer is 28/0.16 =175 fish

28/0.16 = x/1.00